OpenStudy (jordanherron811):

The denominator of a fraction is five more than twice the numerator. If both numerator and denominator are decreased by two, the simplified result is 1 3 . Find the original fraction. (Do NOT simplify.)

OpenStudy (superpanda1022):

I will say x/2x+5

OpenStudy (jordanherron811):

fraction

likeabossssssss (likeabossssssss):

hey welcome to os

OpenStudy (agent0smith):

So you have the original \(\Large \frac{ x }{ 2x+5}\) both numerator and denominator are decreased by two \(\Large \frac{ x -2 }{ 2x+5-2}\) result is 1/3: \(\Large \frac{ x -2 }{ 2x+5-2} =\frac{ 1 }{ 3} \) Solve for x; go back and plug into original fraction.

OpenStudy (jordanherron811):

I don't know how to do that

OpenStudy (agent0smith):

First simplify it, \(\Large \frac{ x -2 }{ 2x+3} =\frac{ 1 }{ 3}\) Then you can cross multiply to start